b1az3:

Find the radical equivalent of 272/3 and compute it.

Xio:

let's find the radical equivalent of 27^(2/3). 27^(2/3) = √27^2 = 3^2 = 9 so your final answer is 9

jhonyy9:

really this 272/3 mean this in this way ?

Xio:

He meant 27^2

jhonyy9:

27^(2/3) mean cube root of 27^2 \[\sqrt[3]{27^{2}} = ?\] \[27 =3^{3}\]

jhonyy9:

hope these will help you

jhonyy9:

\[27^{\frac{ 2 }{ 3}} = \sqrt[3]{27^{2}}\]

jhonyy9:

this is the really problem posted ?

jhonyy9:

@b1az3

jhonyy9:

@pooja195

jhonyy9:

xio do you understand it now the way ?

Xio:

Wot.